Sarah is three more than four times John's age. The difference in their ages is 33. How old is Sarah?
Answer by colliefan(242)

If x=John's age, Sarah's age is 4x+3
If the difference is 33:
(4x+3)-x=33
4x-x+3=33
3x+3=33
3x+3-3=33-3
3x+0=30
3x=30
3x/3=30/3
x=10
If John's age is x, it is 10.
If Sarah's is 4x+3, it is 43.
Is the difference 33? YES!